OpenStudy (anonymous):

Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1).

OpenStudy (anonymous):

start with \(a(x+9)(x+3)\) and then solve for \(a\)

OpenStudy (anonymous):

or start with \(f(x)=a(x+6)^2-1\) and solve for \(a\) either way. you need the steps?

OpenStudy (anonymous):

@Traxter ?

OpenStudy (anonymous):

Ok, so you have y= a(x+9)(x+3) and you need to find a. If we put in x=-6, y=-1 we get: \[-1=a(3)(-3)\] \[-1=-9a\] \[a=\frac{1}{9}\] So we have \[y=\frac{1}{9}(x+9)(x+3)\]

OpenStudy (anonymous):

Please let me know if I can make anything more clear :)

OpenStudy (anonymous):

so that would be the final answer?

OpenStudy (anonymous):

or equation

OpenStudy (anonymous):

@Traxter

OpenStudy (anonymous):

\[y=\frac{1}{9}(x+9)(x+3)\] Or you could expand it: \[y=\frac{1}{9}(x^2+12x+27)\] \[y=\frac{x^2}{9}+\frac{4}{3}x+3\]

OpenStudy (anonymous):

You alright with that?

OpenStudy (anonymous):

yes.thank you :)

OpenStudy (anonymous):

No worries :)